A youth club has 140 members. 80 listen to pop music , 40 to rock music, 75 to heavy metal and 2 members don't listen to any music. 15 members listen to pop and rock only. 12 to rock and heavy metal only, while 10 listen to pop and heavy metal only. How many listen to all three types of music?
Answer by jorel1380(3719)

The club has 140 members, of which 138 listen to music at all. 80 listen to pop,40 listen to rock, 75 to heavy metal. So:
80+40+75=195
And 15+12+10, or 37, listen to 2 types of music.
195-37=158
158-138=20 who listen to all three types of music.
☺☺☺☺